# SPHINX: Structural Prediction using Hypergraph Inference Network

### v1

@@ Abstract
The importance of higher-order relations is widely recognized in a large number of real-world systems. However, annotating them is a tedious and sometimes impossible task. Consequently, current approaches for data modelling either ignore the higher-order interactions altogether or simplify them into pairwise connections. In order to facilitate higher-order processing, even when a hypergraph structure is not available, we introduce Structural Prediction using Hypergraph Inference Network (SPHINX), a model that learns to infer a latent hypergraph structure in an unsupervised way, solely from the final node-level signal. The model consists of a soft, differentiable clustering method used to sequentially predict, for each hyperedge, the probability distribution over the nodes and a sampling algorithm that converts them into an explicit hypergraph structure. We show that the recent advancement in $k$-subset sampling represents a suitable tool for producing discrete hypergraph structures, addressing some of the training instabilities exhibited by prior works. The resulting model can generate the higher-order structure necessary for any modern hypergraph neural network, facilitating the capture of higher-order interaction in domains where annotating them is difficult. Through extensive ablation studies and experiments conducted on two challenging datasets for trajectory prediction, we demonstrate that our model is capable of inferring suitable latent hypergraphs, that are interpretable and enhance the final performance.
@@ 1 Introduction
Graph data is universally recognized as the standard representation for relational data. However, their capabilities are restricted to only modelling pairwise connections. Emerging research shows that real-world applications including neuroscience [1, 2], chemistry [3], biology [4], often exhibit group interactions, involving more that two elements. This leads to the development of a new field dedicated to representing higher-order relations, in the form of hypergraphs.
@@
However, while graph datasets are widespread in the machine learning community [5, 6], the availability of hypergraph datasets is much more limited. Recent work [7] highlights two potential causes for the lack of higher-order data. On one hand, current technology used for collecting information are mostly designed or optimised to detect pairwise interactions. Furthermore, even in the exceptional cases when the data is gathered in a higher-order format, the published version is often released in a reduced, pairwise form.
@@
Therefore, to preserve the higher-order information, it is crucial to develop methods for learning the hypergraph structure in an unsupervised way, only from point-wise observations. Motivated by this, we introduce Structural Prediction using Hypergraph Inference Network (SPHINX), a model for unsupervised latent hypergraph inference, that can be used in conjunction with any recent models designed for hypergraph processing. SPHINX models the hyperedge discovery as a clustering problem, adapting a soft clustering algorithm to sequentially identify subsets of highly-correlated nodes, corresponding to each hyperedge. To produce a discrete hypergraphs structure, we take advantage of the recent development in differentiable $k$-subset sampling [8, 9], obtaining a more effective training procedure, that eliminates the necessity for heavy regularisation. While classical selection methods such as Gumbel-Softmax  [10] fail to control the sparsity of the hypergraph, our constrained $k$-subset sampling produces more accurate latent structure.
@@
Our model improves the relational processing by learning an appropriate hypergraph structure for the downstream task, without any supervision at the structure level. Our experiments on trajectory prediction show that the inferred structure correlates well with the ground-truth connectivity that guides the dynamical process, while also being beneficial for improving the performance on the higher-level task. The resulting model is general, easy to optimise, which makes it an excellent candidate for modelling higher-order relations, even in the absence of an annotated connectivity.
Our main contributions are summarised as follow:
@@
    We propose a novel method for explicit hypergraph inference, that uses a sequential, differentiable clustering approach for identifying subsets of highly-related nodes and a $k$-subset sampling to produce an explicit hypergraph structure, that can be plugged into any hypergraph neural network.
@@
    The model performs unsupervised hypergraph discovery, by using supervision only from the weak node-level signal. We empirically show that the predicted hypergraph correlates well with the true higher-order structure, even if the model was not optimised for this task.
@@
    The latent hypergraph inference enforces an inductive bias towards capturing higher-order correlations, even in the absence of the real structure, which proved to be beneficial for high-level tasks such as trajectory prediction. Having an explicit structure enables us to visualise the discovered hypergraph, adding another layer of interpretability into the model.
@@ 2 Related work
Structural inference on graph.
Modelling relational data using Graph Neural Networks (GNNs) [11, 12] proves to be beneficial in several real-world domains including healthcare [13], physics [14], weather forecast [15], mathematics [16], sports analysis [17], chemistry [18] and many more. Most of the current graph methods assume that the graph connectivity is known. However, providing the relational structure is a highly challenging task, that, when possible to compute, requires either expensive tools, or advanced domain knowledge. These limitations led to the development of a new machine learning field dedicated to learning to infer structure from data in an unsupervised way.
@@
Neural Relational Inference [19] is one of the pioneering works inferring an adjacency matrix from point-wise observations. The model consists of an encoder that learns to predict a distribution over the potential relationships and a GNN decoder that receives a sampled graph structure and learn to predict the future trajectory. fNRI [20] extends this work by inferring a factorised latent space, capable to encode multiple relational types, while  [21] incorporates causal relations into the framework. When using temporal data, these models infer a single structure for the entire timeseries. dNRI [22] improves on that respect, by modifying the graph structure at each timestep.
Hypergraph Networks. While being a versatile, widespread data structure, graphs are constrained to only represent pairwise connections. Many real-world events contains higher-order interactions, where a relationship contains more than two elements. Several data structures for higher-order modelling were proposed in mathematics, including simplicial complexes [23], cell complexes [24], and more generally, hypergraphs [25]. To model hypergraphs, several deep learning methods have recently emerged. Hypergraph Neural Networks [26] applies GNNs on top of a weighted clique expansion. This could be seen as a two-stage message passing scheme. In the first stage the information is send from nodes to hyperedges, while in the second stage the messages are send back from hyperedges to nodes. UniGNN [27] and AllDeepSets [28] introduce a more general framework in which the two stages can be implemented as any permutation-invariant function such as GNNs, DeepSets [29] or Transformers [30]. Further proposed extensions include the integration of attention mechanisms [31, 32, 33], or attaching additional geometric structure to the hypergraph, through the incorporation of cellular sheaves  [34].
Structural inference on hypergraphs.
Providing the hypergraph structure to the model requires measuring higher-order correlations, which often implies a highly difficult and expensive annotation process. This leads to the necessity of inferring the hypergraph structure directly from data. However, when moving from the graph realm to the hypergraph domain, the set of potential edge candidates abruptly increases from quadratic to exponential. This makes the problem of inferring the latent hypergraph structure significantly more challenging. Classical attempts of achieving this includes methods based on Bayesian inference [35] or statistical approaches that compares against a null model for hypergraphs to filter the possible hyperedges [36].
While relational inference for graph data is an established field in deep learning, it remains largely unexplored in the hypergraph domain. Cellular complexes represent an alternative, hierarchical data structure for modelling higher-order relationships. DCM [37] performs structural inference for cellular complexes by proposing a method that learns to efficiently filter from the pool of all possible polygons associated with a graph.  [38] proposes a supervised, iterative refinement method for set-to-hypergraph task, suitable for large input set sizes. For environments when a ground-truth hypergraph structure is not available, most of the existing methods focuses on the transductive setup, where a single hypergraph is inferred for the entire dataset.  [39] treats the incidence matrix as a learnable parameter, while  [40, 41] build a closed-form solution for the label propagation task.
On the other hand, for the inductive setup, where each example requires inferring a distinct hypergraph, the literature is very scarce. GroupNet [42] extracts a subset of nodes by identifying highly connected blocks in the correlation matrix. Alternatively, EvolveHypergraph [43] uses a GNN model to infer the probabilities of each node being part of a hypergraph and transform this into a hypergraph structure using a Gumbel-Softmax [10] sampling technique. While obtaining good results on the trajectory prediction tasks, properly optimising these methods requires a high degree of regularisation techniques, including auxiliary loss terms for reconstruction, smoothness or diversity. We believe that this mainly stems from challenges in effectively propagating gradients during the discretization phase. Gumbel-Softmax struggles to control the cardinality of the hyperedges, while top-$k$ operations only partially tracks gradients. Moreover, these methods suffer a severe drop in performance when using only the hypergraph structure, with best performance requiring additional pairwise computation.
On the other hand, we are proposing a dedicated hypergraph-only model, that can be easily optimised solely from the final, downstream supervision, eliminating the need for additional optimisation heuristics. As we demonstrate in the experimental section, this advantage is brought by two important design choices: 1) the sequential differentiable-clustering which creates globally-informed probability distribution, alleviating the ambiguity existent in the hyperedge prediction problem and 2) adopting the $k$-subset sampling for discretizing the structure, a method that produces high-quality gradients for the constrained sampling operation. The resulting model is a versatile and general module, useful for hypergraph processing in scenarios where the hypergraph is not available.
@@
3 Structure Prediction using Hypergraph Inference Network
@@
The processing stages of our method are summarized in Figure 1. First, we will predict the latent hypergraph based on the input features $\mathbf{X}$. The inferred hypergraph structure will then be fed into a hypergraph network to obtain the final prediction. To obtain a powerful, general model, our method needs to fulfill two important characteristics. Firstly, the entire pipeline need to be end-to-end differentiable, such that the latent hypergraph inference can be trained in a weakly supervised fashion, solely from the final downstream prediction, without access to the ground-truth hypergraph structure. Secondly, the model needs to be general enough such that it can be used inside any existing architecture for hypergraph processing.
To achieve these desiderata, we designed the model using two core components: a learnable soft clustering that predicts a probability distributions for each one of the $M$ potential hyperedges, and a differentiable $k$-sampling that, based on these probability distribution, samples discrete subsets of $k$ nodes forming the hypergraph structure. Both components are differentiable such that the model can be easily trained using standard backpropagation techniques. Moreover, the predicted hypergraph structure is a discrete object, following the classical structural representation used in any hypergraph processing models $\mathcal{H}=(V,E)$, where $V$ is a the of nodes and $E$ is a the of hyperedges.
@@ 3.1 Hypergraph predictor
The goal of this module is to transform a set of node features $\mathbf{X}\in\mathbb{R}^{N\times f}$ into a set of incidence probabilities $\mathbf{P}\in\mathbb{R}^{N\times M}$ where $N$ represents the number of node and $M$ represents the expected number of hyperedges. Each column $j\in\{1\dots M\}$ corresponds to a hyperedge and represents the probability of each node being part of the hyperedge $j$. In order to accurately predict these probabilities, the model needs to have a global understanding of the nodes interactions and identify the subsets that are more likely to exhibit a higher-order relationship.
Taking inspiration from the computer vision literature for unsupervised object detection [44], we model the hypergraph discovery task as a soft clustering problem, where each cluster correspond to a hyperedge. We adapt the iterative slot-attention algorithm [45] to produce $M$ clusters, each one corresponding to a predicted hyperedge. The probability of a node $i$ being part of a cluster $j$, computed as the node-cluster similarity, represents the incidence probabilities $p_{ij}$ corresponding to node $i$ and hyperedge $j$. Therefore, we will use the terms slots and hyperedges interchangeably.
Slot Attention for probabilistic incidence prediction.
We start by creating $M$ slots, one for each hyperedge. Each slot $s_{j}\in\mathbb{R}^{f}$ is randomly initialized from a normal distribution. At each iteration, the slots representation is updated as the weighted average of all the nodes, with the weights computed as a learnable dot-product similarity as indicated by Equation 1, where $f_{1}$, $f_{2}$ and $f_{3}$ represent MLPs and $\sigma$ is a non-linearity. After $Q$ iterations, the pairwise similarity between the updated slot representation and the set of node features represents the predicted probability distribution $p_{:,j}$ for hyperedge $j$.
@@
Note that this procedure is similar to a differentiable $k$-means clusterization, where the hyperedges represent the centroids and the similarity coefficient represent the soft assignment corresponding to each node-cluster pair. At a high level, the probability of a node being part of a hyperedge can be interpreted as the likelihood of an element being assigned to a certain cluster.
Sequential Slot Attention for solving ambiguities.
The above algorithm suffers from a strong limitation. Due to the symmetries exhibited by the set of hyperedges, independently inferring $M$ hyperedges leads to strong ambiguity issues. Concretely, since the slots are initialized randomly, there is no mechanism in place to distribute the hyperedges between slots: multiple slots could be attached to the same obvious hyperedge, leaving others completely uncovered.
To alleviate this, we propose a sequential slot attention. Instead of predicting all hyperedges simultaneously, we will predict them sequentially, one at a time, ensuring that at each timestep the hyperedge prediction mechanism is aware of the hyperedges predicted so far. To achieve that, the features of each node $x_{i}$ is enriched with an additional binary vector $b_{i}\in\{0,1\}^{(M-1)}$ indicating the relationship between that node and the previously predicted hypergraphs. Specifically, when predicting hyperedge $j$, for each previous hyperedge $t$ ($t<j$), $b_{it}=1$ if the node $i$ was previously selected to be part of the hyperedge $t$ and $b_{it}=0$ otherwise. This way, the slot-attention algorithm has the capacity to produce more diverse hyperedges, as we experimentally validated in Section 4.1.

### v2

@@ Abstract
The importance of higher-order relations is widely recognized in numerous real-world systems. However, annotating them is a tedious and sometimes even impossible task. Consequently, current approaches for data modelling either ignore the higher-order interactions altogether or simplify them into pairwise connections. To facilitate higher-order processing, even when a hypergraph structure is not available, we introduce SPHINX, a model that learns to infer a latent hypergraph structure in an unsupervised way, solely from the final task-dependent signal. To ensure broad applicability, we design the model to be end-to-end differentiable, capable of generating a discrete hypergraph structure compatible with any modern hypergraph networks, and easily optimizable without requiring additional regularization losses. Through extensive ablation studies and experiments conducted on four challenging datasets, we demonstrate that our model is capable of inferring suitable latent hypergraphs in both transductive and inductive tasks. Moreover, the inferred latent hypergraphs are interpretable and contribute to enhancing the final performance, outperforming existing methods for hypergraph prediction.
Machine Learning, ICML
@@ 1 Introduction
Graphs are universally recognized as the standard representation for relational data. However, their capabilities are restricted to modelling pairwise connections. Emerging research shows that real-world applications including neuroscience (Guo et al., 2021), chemistry (Jost & Mulas, 2018), biology (Viñas et al., 2022), often exhibit group interactions, involving more than two elements. This leads to the development of a new field dedicated to representing higher-order relations, in the form of hypergraphs. However, while graph datasets are widespread in the machine learning community (Leskovec et al., 2005; Hu et al., 2020), the availability of hypergraph datasets is much more limited. Recent work (Wang & Kleinberg, 2024) highlights two potential causes for the lack of higher-order data. On one hand, current technology used for collecting information is mostly designed or optimised to detect pairwise interactions. Furthermore, even in the exceptional case when the data is gathered in a higher-order format, the published version is often released in a reduced, pairwise form.
Therefore, to preserve the higher-order information, it is crucial to develop methods for learning the hypergraph structure in an unsupervised way, only from point-wise observations. To create a general, widely-applicable model, we identify a set of key desiderata. 1) The model should be applicable to a broad type of tasks. This is a challenge for most existing hypergraph predictors, which optimize a single hypergraph, limiting their usefulness to transductive setups. 2) The model should be compatible with any hypergraph processing architecture. This requires our inference model to be fully differentiable while producing a sparse and discrete hypergraph structure. Existing methods often fall short, as they are either only partially differentiable (relying on techniques such as top-k selection or thresholding to enforce sparsity) or restricted to generating only weighted hypergraphs. 3) A powerful model should be easy to optimise. This is a limitation exhibited by most of the modern architectures for which predicting a suitable hypergraph requires various types of regularization losses.
@@
Motivated by these, we introduce Structural Prediction using Hypergraph Inference Network (SPHINX), a model for unsupervised latent hypergraph inference, that can be used in conjunction with any recent models designed for hypergraph processing. SPHINX models the hyperedge discovery as a clustering problem, adapting a soft clustering algorithm to sequentially identify subsets of highly correlated nodes, corresponding to each hyperedge. To produce a discrete hypergraph structure, we take advantage of the recent development in differentiable $k$-subset sampling (Ahmed et al., 2023; Minervini et al., 2023), obtaining a more effective training procedure, that eliminates the necessity for heavy regularisation. While classical selection methods such as Gumbel-Softmax  (Jang et al., 2017) fail to control the sparsity of the hypergraph, our constrained $k$-subset sampling produces a more accurate latent structure.
@@
Experiments on both inductive and transductive datasets show that our inferred hypergraph surpasses existing graph and hypergraph-based models on the final downstream task. Moreover, our synthetic experiments reveal that the inferred structure correlates well with the ground-truth connectivity that guides the dynamical process, outperforming other hypergraph predictors. The resulting model is general, easy to optimise, which makes it an excellent candidate for modelling higher-order relations, even in the absence of an annotated connectivity.
Our main contributions are summarised as follows:
@@
    We propose a novel method for explicit hypergraph inference, that uses a sequential predictor to identify subsets of highly-related nodes and a $k$-subset sampling to produce an explicit hypergraph structure, that can be plugged into any hypergraph neural network.
@@
    The model performs unsupervised hypergraph discovery, by using supervision only from the weak node-level signal. We empirically show that the predicted hypergraph correlates well with the true higher-order structure, even if the model was not optimised for it.
@@
    The latent hypergraph enforces an inductive bias for capturing higher-order correlations, even in the absence of the real structure, which proved to be beneficial for downstream tasks such as trajectory prediction (inductive) or node classification (transductive). Having an explicit structure allows us to visualise the discovered hypergraph, adding a new layer of interpretability to the model.
@@ 2 Related work
Structural inference on graphs. Modelling relational data using Graph Neural Networks (GNNs) (Kipf & Welling, 2017; Veličković et al., 2018) proves to be beneficial in several real-world domains (Bica & van der Schaar, 2022; Sanchez-Gonzalez et al., 2020; Lam et al., 2023). Most of the current graph methods assume that the graph connectivity is known. However, providing the relational structure is a highly challenging task, that, when possible to compute, requires either expensive tools, or advanced domain knowledge. These limitations led to the development of a new machine learning field dedicated to learning to infer structure from data in an unsupervised way. Neural Relational Inference (Fetaya et al., 2018) is one of the pioneering works inferring an adjacency matrix from point-wise observations. The model consists of an encoder that learns to predict a distribution over the potential relationships and a GNN decoder that receives a sampled graph structure and learns to predict the future trajectory. fNRI (Webb et al., 2019) extends this work by inferring a factorised latent space, capable of encoding multiple relational types, while  (Lowe et al., 2022) incorporates causal relations into the framework. For temporal data, these models infer a single structure for the entire timeseries. dNRI (Graber & Schwing, 2020) improves on that respect, by modifying the graph structure at each timestep.
@@
Hypergraph Networks. Several data structures for higher-order modelling were proposed, including simplicial complexes (Torres & Bianconi, 2020), cell complexes (Lundell & Weingram, 1969), and more generally, hypergraphs (Banerjee, 2021). To model hypergraphs, several deep learning methods have recently emerged. Hypergraph Neural Networks (Feng et al., 2019) applies GNNs on top of a weighted clique expansion. This could be seen as a two-stage message passing scheme, sending messages from nodes to hyperedges and viceversa. UniGNN (Huang & Yang, 2021) and AllDeepSets (Chien et al., 2022) introduce a more general framework in which the two stages can be implemented as any permutation-invariant function such as DeepSets (Zaheer et al., 2017) or Transformers (Vaswani et al., 2017). Further proposed extensions include the integration of attention mechanisms (Bai et al., 2021; Zhang et al., 2022; Georgiev et al., 2022), or attaching additional geometric structure to the hypergraph, through the incorporation of cellular sheaves  (Duta et al., 2023). While these methods can be seen as learning to modify the structure, they require an initial hypergraph structure.
Structural inference on hypergraphs. Providing the hypergraph structure to the model requires measuring higher-order correlations, which often implies a highly difficult and expensive annotation process. This leads to the necessity of inferring the latent hypergraph structure directly from data. However, when moving from the graph realm to the hypergraph domain, the set of potential edge candidates abruptly increases from quadratic to exponential. This makes the problem of inferring the latent hypergraph structure significantly more challenging. Classical attempts of achieving this include methods based on Bayesian inference (Young et al., 2021) or statistical approaches that filter against a null model for hypergraphs (Musciotto et al., 2021).
While relational inference for graphs is an established field in deep learning, it remains largely underexplored in the hypergraph domain. The common approach is to either extract hyperedges using k-NN (Huang et al., 2009), clustering (Jiang et al., 2019; Gao et al., 2013) or iteratively optimising regularisation constraints (Liu et al., 2016). However, the resulting hypergraph is static, independent of the targeted downstream task. More recent several methods (Yu et al., 2012; Wang et al., 2024; Cai et al., 2022; Zhang et al., 2018; Gao et al., 2020) improve on the kNN approach by learning a task-guided mask to modify the initial structure. While producing dynamic hypergraphs, optimising the mask restrict the method to only work in the transductive setup, in which a single hypergraph is produced for the entire dataset.
Very few methods in the literature are general enough to allow learning hypergraph structure in both inductive and transductive tasks. [^1] From now on we use the term inductive when a different hypergraph is predicted for each example and transductive setup when a single hypergraph is predicted for the entire dataset. GroupNet (Xu et al., 2022) and DynGroupNet (Xu et al., 2024) use the correlation matrix to extract submatrices of high connectivity, EvolveHypergraph (Li et al., 2022) predicts a soft connectivity for each node, relying on Gumbel-Softmax and regularisation tricks to produce a sparse, discrete structure, while TDHNN (Zhou et al., 2023) uses a differentiable clustering to produce soft incidence, followed by a top-k selection to impose sparsity. For a more in-depth analysis of the existing dynamic hypergraph predictors, please see Section B of the Appendix.
3 Structure Prediction using SPHINX
@@
The processing stages of our method are summarized in Figure 1. First, we predict the latent hypergraph based on the input features $\mathbf{X}$. The inferred hypergraph structure is then fed into a hypergraph network to obtain the final prediction. To fulfil our goal of obtaining a generally applicable method, easy to optimize and compatible with any hypergraph architecture, our method needs to satisfy two important characteristics. Firstly, the pipeline needs to be end-to-end differentiable, such that the latent hypergraph inference can be trained in a weakly supervised fashion, solely from the downstream signal. Secondly, it needs to produce a sparse and discrete structure, such that it can be used inside any existing hypergraph processing model.
To achieve these, we designed the model using two core components: a learnable soft clustering that sequentially predicts a probability distributions for each one of the $M$ potential hyperedges, and a differentiable $k$-sampling that, based on this probability distribution, samples discrete subsets of $k$ nodes forming the hypergraph structure. Both components are differentiable such that the model can be easily trained using standard backpropagation techniques. Moreover, the predicted hypergraph structure is a discrete object, following the classical structural representation used in any hypergraph processing models $\mathcal{H}=(V,E)$, where $V$ is a the of nodes and $E$ is a the of hyperedges.
@@ 3.1 Hypergraph predictor
The goal of this module is to transform a set of node features $\mathbf{X}\in\mathbb{R}^{N\times f}$ into a set of incidence probabilities $\mathbf{P}\in\mathbb{R}^{N\times M}$ where $N$ represents the number of nodes and $M$ represents the expected number of hyperedges. Each column $j\in\{1\dots M\}$ corresponds to a hyperedge and represents the probability of each node being part of the hyperedge $j$. In order to accurately predict these probabilities, the model needs to have a global understanding of the nodes interactions and identify the subsets that are more likely to exhibit a higher-order relationship.
Taking inspiration from the computer vision literature for unsupervised object detection (Greff et al., 2020), we model the hypergraph discovery task as a soft clustering problem, where each cluster corresponds to a hyperedge. We adapt the iterative slot-attention algorithm (Locatello et al., 2020) to produce $M$ clusters, each one corresponding to a predicted hyperedge. The probability of a node $i$ being part of a cluster $j$, computed as the node-cluster similarity, represents the incidence probabilities $p_{ij}$ corresponding to node $i$ and hyperedge $j$. Therefore, we will use the terms slots and hyperedges interchangeably.
Slot Attention for probabilistic incidence prediction. We start by creating $M$ slots, one for each hyperedge. Each slot $s_{j}\in\mathbb{R}^{f}$ is randomly initialized from a normal distribution. At each iteration, the slots representation is updated as the weighted average of all the nodes, with the weights computed as a learnable dot-product similarity as indicated by Equation 1, where $f_{1}$, $f_{2}$ and $f_{3}$ represent MLPs and $\sigma$ is a non-linearity. After $Q$ iterations, the pairwise similarity between the updated slot representation and the set of node features represents the predicted probability distribution $p_{:,j}$ for hyperedge $j$.
@@
Sequential Slot Attention for solving ambiguities. The above algorithm suffers from a strong limitation. Due to the symmetries exhibited by the set of hyperedges, independently inferring $M$ hyperedges leads to strong ambiguity issues. Concretely, since the slots are initialized randomly, there is no mechanism in place to distribute the hyperedges between slots: multiple slots could be attached to the same obvious hyperedge, leaving others completely uncovered.
To alleviate this, we propose a sequential slot attention. Instead of predicting all hyperedges simultaneously, we will predict them sequentially, one at a time, ensuring that at each timestep the hyperedge prediction mechanism is aware of the hyperedges predicted so far. To achieve that, the features of each node $x_{i}$ are enriched with an additional binary vector $b_{i}\in\{0,1\}^{(M-1)}$ indicating the relationship between that node and the previously predicted hypergraphs. Specifically, when predicting hyperedge $j$, for each previous hyperedge $t$ ($t<j$), $b_{it}=1$ if the node $i$ was previously selected to be part of the hyperedge $t$ and $b_{it}=0$ otherwise. This way, the slot-attention algorithm has the capacity to produce more diverse hyperedges, as we show in Section 4.1.